OpenStudy (anonymous):

what is -3/276(36x10)=?

OpenStudy (thomas5267):

A or B or neither? \[ \begin{align*} &\frac{-3}{276}(36\times10)&\text{(a)}\\ &\frac{-3}{276(36\times10)}&\text{(b)} \end{align*} \]

OpenStudy (anonymous):

A

OpenStudy (thomas5267):

Hint: \(276=2\times 2\times 23\).

OpenStudy (anonymous):

I keep getting 92=276 and idk wat I do wrong

OpenStudy (thomas5267):

How do you get 92=276?

OpenStudy (anonymous):

I did 2x2x23 first

OpenStudy (thomas5267):

276=2x2x3x23 I am very sorry... :(

OpenStudy (anonymous):

ooooh ok thank u and thts ok we all make mistakes